Mathematics 242 – Principles of Analysis Definitions Quiz 6 – November 5, 2004

Your Name:

1. (15) Give the precise definition of the statement a = sup(S) for a subset S of the real numbers

2. (10) What does the Least Upper Bound (Completeness) Axiom for the real number system say?